# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 3.4. Properties of the kernelized instances
Note that a sharp subtree of $T$ contains a path of size at least $2$ and whose endpoints are adjacent to a same vertex $u\in M$. So as rule (KR5) does not apply, we have that $T$ contains at most $d_{M}(T)(d_{M}(T)+b_{\overline{M}}(T))$ disjoint sharp subtrees.
Let $T^{1}=T$ (and remember that the trees of $G-M$ are rooted). Then, iteratively let $v_{i}$ be a lowest vertex of $T^{i}$ such that $T^{i}_{v_{i}}$ is sharp, and let $T^{i+1}=T^{i}\setminus T^{i}_{v_{i}}$. This process stops when the remaining tree is weakly connected to $M$ or empty. This leads to a partition of $T$ into $c+1$ disjoint subtrees, where the first $c$ are sharp subtrees, and the last one is weakly connected to $M$, with $c\leq d_{M}(T)(d_{M}(T)+b_{\overline{M}}(T))=\mathcal{O}(\operatorname{\mu}(T)^{2})$. Using the bound on the size of the weakly connected and sharp trees from 3.16 and 3.17, we get:
@@
Given an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2), and (KR4) applies, and a subtree $T$ of $G-M$ with $|T|>p_{1}(r,\operatorname{\mu}(T))$, Rule (KR5) can be applied in polynomial time on this tree.
@@
The fact that Rule (KR5) applies is the contrapositive of 3.19, and the $d_{M}(T)+b_{\overline{M}}(T)$ paths needed to apply the rule can be found using the partition used in its proof, which can be computed in polynomial time. ∎
@@ 3.6. Branching when there is a $\widetilde{K_{t,t}}$
When reaching line 5 of Algorithm 1, we call the method Partitionning-Algorithm  which tries to build a special partition of $G-M$. As we will see later, one output of this procedure is a failure (called case 2) where a certain dense structure (denoted $\widetilde{K_{t,t}}$, with $t=2d_{r}$) is found. In this section, we explain how we branch on such a $\widetilde{K_{t,t}}$.
This new branching rule is a variation of Rule (BR1) that was presented in Subsection 3.1: instead of dealing with $K_{r,r}$ subgraphs, we will now consider a set $A$ of $t$ vertices and a set $B$ of small trees such that contracting those trees would result in a $K_{t,t}$-subgraph. More formally:
@@
Given an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS and an integer $t$, a $\widetilde{K_{t,t}}$ (of $(G,k,M,\mathcal{H})$) is a pair $\left(X,(T_{i})_{1\leq i\leq t}\right)$ where
@@
  • $(T_{i})_{i}$ a family of $t$ disjoint non-adjacent subtrees of $G-M$ such that for all $1\leq i\leq t$, $X\subseteq N_{M}(T_{i})$, and
  • for any $T_{i}$, $|T_{i}|\leq p_{2}(r,t)$ with $p_{2}(r,t)=p_{1}(r,2t)$.
@@
Similarly to a $K_{t,t}$ subgraph, given a $\widetilde{K_{t,t}}$ with parts $(X,(T_{i})_{1\leq i\leq t})$, the subgraph of $G$ induced by two vertices of $X$ and two trees of $(T_{i})_{i}$ always contains a cycle. So a feedback vertex set of $G$ must either contain at least $t-1$ vertices of $X$, or intersects each tree of $(T_{i})_{i}$ except at most one.
The difference between the treatment of $K_{t,t}$ and $\widetilde{K_{t,t}}$ is that, when we are in the branch where a solution must intersect each tree of $(T_{i})_{i}$, we cannot branch in subexponential time to guess which vertex of each $T_{i}$ is picked, and rather perform the following “virtual branching” trick (introduced in [LPS+22] for the special case where $T_{i}$ are edges). In this branch, we add all these $T_{i}$ in $M$, and maintain a packing $\mathcal{H}$ of such sets $T_{i}$. The size of the packing (and thus the growth of $M$) will be bounded by observing that $|\mathcal{H}|$ is a lower bound on the solution size.
We are now ready to define the last branching rule:
@@
    Given an instance $(G,k,M,\mathcal{H})$ and a $\widetilde{K_{t,t}}$ with parts $(X,(T_{i})_{i})$, we generate the following instances:
@@
    • if $|\mathcal{H}|+(t-1)\leq k$, then for each $1\leq i\leq t$, denoting $R_{i}=\bigcup_{j\neq i}T_{j}$ and $\mathcal{T}_{R_{i}}=\{T_{j},~{}j\neq i\}$, output the instance $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$.
@@
The Rule (BR2) is safe, and can be applied in polynomial time (assuming a $\widetilde{K_{t,t}}$ is provided).
@@
The obtained graph is an induced subgraph of the original graph, so the rule preserves the property of being $K_{r,r}$-free $s$-string graph. The running time bound is straightforward.
The fact that the input and output instances are equivalent comes from 3.22. More formally, suppose that $S$ is a solution to $(G,k,M,\mathcal{H})$. If $|S\cap X|\geq t-1$, then there exists $v\in X$ such that $S\supseteq X_{\overline{v}}$, implying that $((G-X_{\overline{v}}),k-(t-1),M\setminus X_{\overline{v}},\mathcal{H}-X_{\overline{v}})$ is a yes-instance. Otherwise ($|S\cap X|\leq t-2$), there exists $u,v$ two vertices in $X\setminus S$. As for any $i$, $G[T_{i}]$ is a tree and $X\subseteq N_{M}(T_{i})$, there cannot be $i\neq j$ such that $S\cap T_{i}=S\cap T_{j}=\emptyset$, as otherwise $G[\{u,v\}\cup T_{i}\cup T_{j}]$ would contain a cycle not hit by $S$. This implies that there exists $i$ such that all trees of $\mathcal{T}_{R_{i}}$ are hit by $S$. As $S$ must also hit any tree in $\mathcal{H}$ (which are disjoint), it implies that $|\mathcal{H}|+t-1\leq k$, and thus the rule generates in particular $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$, which is a YES-instance. The reverse direction of safeness is immediate.
3.7. Attempting to build a $t$-uniform partition of $G-M$
In this section we define the method Partitionning-Algorithm  that is called line 5 when (KR1), (KR2), (KR3), and (KR4) do not apply. The algorithm will either fail (case 1) and find a “big” subtree $T$ that allows to apply (KR3) of Subsection 3.5, fail (case 2) and find a $\widetilde{K_{t,t}}$ that allows to apply (BR2) of Subsection 3.6, or find a special partition of $G-M$ that we define now.
@@ 3.7. Attempting to build a $t$-uniform partition of $G-M$
Let $(G,k,M,\mathcal{T})$ be an instance of $(r,\mathcal{G})$-Ann-FVS and $t$ be a positive integer. Let $F\subseteq V(G)\setminus M$ be a downward-closed subforest of $G-M$. A family of trees $\mathcal{T}$ of $G-M$ with vertices in $F$ is a $t$-uniform partition of $F$ if:
@@
    Each vertex of $F$ is in exactly one tree of $\mathcal{T}$.
@@
    Every $T\in\mathcal{T}$ have $d_{M}(T)\leq 2t$, $|T|\leq p_{1}(r,\operatorname{\mu}(T))$ (where $p_{1}$ is the polynomial function defined in 3.19) and $\mathcal{T}$ is partitioned into two subsets, $\mathcal{T}^{-}$ and $\mathcal{T}^{+}$ such that a tree $T\in\mathcal{T}$ belongs to $\mathcal{T}^{-}$ if $d_{M}(T)<t$, and to $\mathcal{T}^{+}$ otherwise.
@@
    If a root $r$ of a tree $\mathcal{T}^{-}$ has a parent $v$ in $G-M$, then $v$ is the root of a tree in $\mathcal{T}^{+}$.
@@
The definition implies that there is no edges between trees of $\mathcal{T}^{-}$.
Figure 4. Representation of a $t$-uniform partition of a subtree $T$ of $G-M$. Here are represented only vertices of $F$ (and not vertices of $M$). Vertices of the trees in $\mathcal{T}^{+}$ (respectively $\mathcal{T}^{-}$) are represented in red (respectively blue), so as the edges between two vertices in the same tree of $\mathcal{T}^{+}$ (respectively $\mathcal{T}^{-}$). Edges between distinct trees of $\mathcal{T}$ are represented in black.
Our goal is to apply the branching rule each time the procedure of 3.27 end up in the item 2, giving a large number of saturated trees. We prove that it can always be done, and even in polynomial time:
When Partitionning-Algorithm  tries to build a $t$-uniform partition, one case (case 2) of failure will be when $\mathcal{T}^{+}$ becomes to large, and thus we show in the next lemma that in this case we can find a $\widetilde{K_{t,t}}$.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2), and (KR4) applies, a downward-closed subforest $F\subseteq V(G-M)$ and a $t$-uniform partition $\mathcal{T}$ of $F$ with $|\mathcal{T}^{+}|=p_{3}(r,t)|M|$ for $p_{3}(r,t)=4tf_{2}(r,2t,p_{2}(r,t))$. Then, we can find a $\widetilde{K_{t,t}}$ in polynomial time.
@@
Consider $F_{\mathcal{T}^{+}}$ the forest obtained from $G-M$ by contracting each subtree in $\mathcal{T}$ to a single vertex, and removing the vertices of the trees in $\mathcal{T}^{-}$. Denoting $V_{s}$ the set of vertices of $F_{\mathcal{T}^{+}}$ with degree at most $2$, we have $|V_{s}|\geq|\mathcal{T}^{+}|/2$ as the number of leaves in a tree is an upper bound on the number of vertices of degree bigger than $3$. Then using a 2-coloration of the forest we partition the vertices of $F_{\mathcal{T}^{+}}$ in two independent sets, $V_{1}$ and $V_{2}$. Without loss of generality we suppose that $|V_{s}\cap V_{1}|\geq|V_{s}\cap V_{2}|$. So denoting $\mathcal{T}^{\prime}$ the set of trees in $\mathcal{T}^{+}$ associated to the vertices in $V_{s}\cap V_{1}$, we have $|\mathcal{T}^{\prime}|\geq|V_{s}|/2\geq|\mathcal{T}^{+}|/4$. Now for $T\in\mathcal{T}^{\prime}$ we want to bound the size of its border. Firstly, for the trees in $\mathcal{T}^{+}$ adjacent to $T$, there are at most two of them by construction of $\mathcal{T}^{\prime}$. Secondly, by item 3 of 3.24 the only vertex in $T$ which can be adjacent to a tree $T^{\prime}\in\mathcal{T}^{-}$ is the root of $T$. So overall, the border of $T$ contains at most $3$ vertices, and so $b_{\overline{M}}(T)\leq 3$. Moreover, as $V_{1}$ is an independent set, there is no edge between two trees of $\mathcal{T}^{\prime}$. By 3.24, any tree of $T\in\mathcal{T}^{\prime}$ has size at most $p_{1}(r,\operatorname{\mu}(T))$, but we have $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))=d_{M}(T)\leq 2t$ so $|T|\leq p_{1}(r,2t)=p_{2}(r,t)$. Let us now apply the second result of 3.13 with $x=t$, $A=M$, $p=2t$ and $m=p_{2}(r,t)$. As $|\mathcal{T}^{\prime}|\geq|\mathcal{T}^{+}|/4=tf_{2}(r,2t,p_{2}(r,t))|M|$, there exists a subset $X\subseteq M$ and a subset $\mathcal{T}^{\prime}_{X}\subseteq\mathcal{T}^{\prime}$ such that $|\mathcal{T}^{\prime}_{X}|=t$, and for all $T\in\mathcal{T}^{\prime}_{X}$, we have $N_{M}(T)=X$. Observe that as $\mathcal{T}^{\prime}\subseteq\mathcal{T}^{+}$, we have $|X|\geq t$, so we can take any subset $X_{0}$ of $X$ such that $|X_{0}|=t$ and then $(X_{0},\mathcal{T}^{\prime}_{X})$ is a $\widetilde{K_{t,t}}$, as wanted.

### v2

@@ 3.4. Properties of the kernelized instances
Note that a sharp subtree of $T$ contains a path whose endpoints are distinct and adjacent to a same vertex $u\in M$. So as the rule (KR5) does not apply, we have that $T$ contains at most $d_{M}(T)(d_{M}(T)+b_{\overline{M}}(T))$ disjoint sharp subtrees.
Let $T^{1}=T$ (and remember that the trees of GMG-M are rooted). Then, iteratively let $v_{i}$ be a lowest vertex of $T^{i}$ such that $T^{i}_{v_{i}}$ is sharp, and let $T^{i+1}=T^{i}\setminus T^{i}_{v_{i}}$. This process stops when the remaining tree is weakly connected to $M$ or empty. This leads to a partition of $T$ into $c+1$ disjoint subtrees, where the first $c$ are sharp subtrees, and the last one is weakly connected to $M$, with $c\leq d_{M}(T)(d_{M}(T)+b_{\overline{M}}(T))=\mathcal{O}(\operatorname{\mu}(T)^{2})$. Using the bound on the size of the weakly connected and sharp trees from 3.14 and 3.15, we get:
@@
Given an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2), and (KR4) applies, and given a subtree $T$ of GMG-M with $|T|>p_{1}(r,\operatorname{\mu}(T))$, Rule (KR5) can be applied in polynomial time on this tree.
@@
The fact that Rule (KR5) applies is the contrapositive of 3.17, and the $d_{M}(T)+b_{\overline{M}}(T)$ paths needed to apply the rule can be found using the partition used in its proof, which can be computed in polynomial time. ∎
@@ 3.6. Branching when there is a $\widetilde{K_{t,t}}$
When reaching line 5 of Algorithm 1, we call the method Partitionning-Algorithm  which tries to build a special partition of GMG-M. As we will see later, one of the possible outputs of this procedure is a failure (called Case 2) where a certain dense structure called $\widetilde{K_{t,t}}$ is found (for $t=2d(r)$). In this section, we explain how we handle such a $\widetilde{K_{t,t}}$. Informally, a $\widetilde{K_{t,t}}$ in an instance $(G,k,M\mathcal{H})$ is a special type of minor model of $K_{t,t}$ where the vertices of one side of the bipartition of $K_{t,t}$ correspond to single vertices of $M$ and the vertices of the other side of the bipartition correspond to small trees of GMG-M. It is formally defined as follows.
@@
For an integer $t$, a $\widetilde{K_{t,t}}$ in an instance $(G,k,M,\mathcal{H})$ of $(r,\mathcal{G})$-Ann-FVS is a pair $\left(X,\{T_{i}\}_{1\leq i\leq t}\right)$ where
@@
  • $\{T_{i}\}_{1\leq i\leq t}$ a family of $t$ disjoint non-adjacent subtrees of GMG-M such that for all $1\leq i\leq t$, $X\subseteq N_{M}(T_{i})$, and
  • for every $i\in\left\{1,\dots,t\right\}$, $|T_{i}|\leq p_{2}(r,t)$ with $p_{2}(r,t)=p_{1}(r,2t)$, where $p_{1}$ is the polynomial defined in 3.17.
@@
Similarly to a $K_{t,t}$ subgraph, given $(X,\{T_{i}\}_{1\leq i\leq t})$ a $\widetilde{K_{t,t}}$, any subgraph of $G$ induced by two vertices of $X$ and two trees of $\{T_{i}\}_{1\leq i\leq t}$ contains a cycle. So a feedback vertex set of $G$ either contains at least t1t-1 vertices of $X$, or it intersects all but at most one of the trees of $\{T_{i}\}_{1\leq i\leq t}$.
This new branching rule is a variation of Rule (BR1) to deal with $\widetilde{K_{t,t}}$’s instead of $K_{t,t}$ subgraphs. The difference between the treatment of $K_{t,t}$ and $\widetilde{K_{t,t}}$ is that, when we are handling the case where a solution must intersect each tree of $\{T_{i}\}_{1\leq i\leq t}$, we cannot branch in subexponential time to guess which vertex of each $T_{i}$ is picked, and rather perform the following “virtual branching” trick that generalizes that of [lokSODA22] corresponding to the special case where $T_{i}$ are edges. In this branch, we add all these $T_{i}$’s in $M$, and add the vertex sets of each tree to $\mathcal{H}$. The size of $\mathcal{H}$ (and thus the growth of $M$) will be bounded by observing that since the sets in $\mathcal{H}$ are disjoint, $|\mathcal{H}|$ is a lower bound on the solution size.
We are now ready to define our last branching rule:
@@
    Given $(G,k,M,\mathcal{H})$ an instance of $(r,\mathcal{G})$-Ann-FVS and $(X,\{T_{i}\}_{i})$ a $\widetilde{K_{t,t}}$, we branch on the following (at most $2t$) instances:
@@
    • if $k\geq|\mathcal{H}|+t-1$, then for each $1\leq i\leq t$, denoting $R_{i}=\bigcup_{j\neq i}T_{j}$ and $\mathcal{T}_{R_{i}}=\{T_{j},~j\neq i\}$, output the instance $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$.
    • if none of the above applies, output a trivial No-instance.
@@
Given $(G,k,M,\mathcal{H})$ an instance of $(r,\mathcal{G})$-Ann-FVS and $(X,\{T_{i}\}_{i})$ a $\widetilde{K_{t,t}}$, the rule (BR2) produces a set of instances of $(r,\mathcal{G})$-Ann-FVS such that $(G,k,M,\mathcal{H})$ is a positive instance if and only if one of the output instances is positive. Moreover, the rule can be performed in polynomial time.
@@
The obtained graph is an induced subgraph of the original graph, so it belongs to $\mathcal{G}$ and is $K_{r,r}$-free. Also, recall that $\mathcal{H}$ consists of disjoint subsets of $M$. In particular, in the second item of (BR2) these sets are disjoint from those in $\mathcal{T}_{R_{i}}$ as for every $i\in\left\{1,\dots,t\right\}$, $T_{i}$ is a subtree in GMG-M. Therefore $\mathcal{H}\cup\mathcal{T}_{R_{i}}$ is a collection of disjoint subsets of $M\cup R_{i}$. Hence the rule (BR2) indeed outputs instances of $(r,\mathcal{G})$-Ann-FVS. The running time bound is straightforward as it is not required to find the $\widetilde{K_{t,t}}$.
The fact that the input and output instances are equivalent comes from 3.20. More formally, suppose that $S$ is a solution to $(G,k,M,\mathcal{H})$. If $|S\cap X|\geq t-1$, then there exists $v\in X$ such that $S\supseteq X_{\overline{v}}$, implying that $((G-X_{\overline{v}}),k-(t-1),M\setminus X_{\overline{v}},\mathcal{H}-X_{\overline{v}})$ is a Yes-instance. Otherwise ($|S\cap X|\leq t-2$), there exists $u,v$ two vertices in $X\setminus S$. As for any $i$, $G[T_{i}]$ is a tree and $X\subseteq N_{M}(T_{i})$, there cannot be $i\neq j$ such that $S\cap T_{i}=S\cap T_{j}=\emptyset$, as otherwise $G[\{u,v\}\cup T_{i}\cup T_{j}]$ would contain a cycle not hit by $S$. This implies that there exists $i$ such that all trees of $\mathcal{T}_{R_{i}}$ are hit by $S$. As $S$ must also hit any tree in $\mathcal{H}$ (which are disjoint), it implies that $k\geq|\mathcal{H}|+t-1$, and thus the rule generates in particular $\left(G,k,M\cup R_{i},\mathcal{H}\cup\mathcal{T}_{R_{i}}\right)$, which is a Yes-instance. The reverse direction is immediate. ∎
3.7. Attempting to build a $t$-uniform partition of GMG-M
In this section we define the routine Partitionning-Algorithm  that is called line 5 when (KR1), (KR2), (KR3), and (KR4) do not apply. The algorithm will either fail (Case 1) and find a large enough subtree $T$ that allows to apply (KR5) of Subsection 3.5, fail (Case 2) and find a $\widetilde{K_{t,t}}$ that allows to apply (BR2) of Subsection 3.6, or find a certain partition of GMG-M that we define now.
@@ 3.7. Attempting to build a $t$-uniform partition of GMG-M
Let $(G,k,M,\mathcal{H})$ be an instance of $(r,\mathcal{G})$-Ann-FVS and $t$ be a positive integer. Let $F\subseteq V(G-M)$ be a downward-closed subforest of GMG-M. A $t$-uniform partition of $F$ is family $\mathcal{T}$ of subtrees of $F$ such that:
@@
    each vertex of $F$ is in exactly one tree of $\mathcal{T}$,
@@
    for every $T\in\mathcal{T}$ have $d_{M}(T)\leq 2t$ and $|T|\leq p_{1}(r,\operatorname{\mu}(T))$ (where $p_{1}$ is the polynomial function defined in 3.17), and
@@
    if a root $r$ of a tree of $\mathcal{T}^{-}$ has a parent $v$ in GMG-M, then $v$ is the root of a tree in $\mathcal{T}^{+}$,
where $\mathcal{T}^{-}$ denote the subset of those $T\in\mathcal{T}$ such that $d_{M}(T)<t$ and $\mathcal{T}^{+}$ the others.
@@
The definition above implies that there are no edges between the trees of $\mathcal{T}^{-}$.
Figure 4. Representation of a $t$-uniform partition of a subtree $T$ of GMG-M. Here are represented only vertices of $F$ (and not vertices of $M$). Vertices of the trees in $\mathcal{T}^{+}$ (respectively $\mathcal{T}^{-}$) are represented in red (respectively blue), so as the edges between two vertices in the same tree of $\mathcal{T}^{+}$ (respectively $\mathcal{T}^{-}$). Edges between distinct trees of $\mathcal{T}$ are represented in black.
One case of failure of Partitionning-Algorithm  when trying to build a $t$-uniform partition is when $\mathcal{T}^{+}$ becomes too large. We show in the next lemma that in this case we can find a $\widetilde{K_{t,t}}$.
@@
Consider an instance $(G,k,M,\mathcal{H})$ where none of Rules (KR1), (KR2), and (KR4) applies, a downward-closed subforest $F\subseteq V(G-M)$ and a $t$-uniform partition $\mathcal{T}$ of $F$ with $|\mathcal{T}^{+}|\geq p_{3}(r,t)|M|$ for $p_{3}(r,t)=4tf_{2}(r,2t,p_{2}(r,t))$. Then, we can find a $\widetilde{K_{t,t}}$ in polynomial time.
@@
Consider $F_{\mathcal{T}^{+}}$ the forest obtained from GMG-M by contracting each subtree in $\mathcal{T}^{+}$ to a single vertex, and removing the vertices of the trees in $\mathcal{T}^{-}$. Denoting $V_{s}$ the set of vertices of $F_{\mathcal{T}^{+}}$ with degree at most $2$ in $F_{\mathcal{T}^{+}}$. We have $|V_{s}|\geq|\mathcal{T}^{+}|/2$ as the number of leaves in a tree is an upper bound on the number of vertices of degree at least $3$. Then using a 2-coloration of the forest we partition the vertices of $F_{\mathcal{T}^{+}}$ in two independent sets, $V_{1}$ and $V_{2}$. Without loss of generality we suppose that $|V_{s}\cap V_{1}|\geq|V_{s}\cap V_{2}|$. So denoting $\mathcal{T}^{\prime}$ the set of trees in $\mathcal{T}^{+}$ associated to the vertices in $V_{s}\cap V_{1}$, we have $|\mathcal{T}^{\prime}|\geq|V_{s}|/2\geq|\mathcal{T}^{+}|/4$. Now for $T\in\mathcal{T}^{\prime}$ we want to bound the size of its border. Firstly, for the trees in $\mathcal{T}^{+}$ adjacent to $T$, there are at most two of them by construction of $\mathcal{T}^{\prime}$. Secondly, by Item 3 of 3.22 the only vertex in $T$ which can be adjacent to a tree $T^{\prime}\in\mathcal{T}^{-}$ is the root of $T$. So overall, the border of $T$ contains at most $3$ vertices, and so $b_{\overline{M}}(T)\leq 3$. Moreover, as $V_{1}$ is an independent set, there is no edge between two trees of $\mathcal{T}^{\prime}$. By 3.22, any tree of $T\in\mathcal{T}^{\prime}$ has size at most $p_{1}(r,\operatorname{\mu}(T))$, but we have $\operatorname{\mu}(T)=\max(d_{M}(T),b_{\overline{M}}(T))=d_{M}(T)\leq 2t$ so $|T|\leq p_{1}(r,2t)=p_{2}(r,t)$. Let us now apply the second result of 3.11 with $x=t$, $A=M$, $p=2t$ and $m=p_{2}(r,t)$. As $|\mathcal{T}^{\prime}|\geq|\mathcal{T}^{+}|/4\geq tf_{2}(r,2t,p_{2}(r,t))|M|$, there exists a subset $X\subseteq M$ and a subset $\mathcal{T}^{\prime}_{X}\subseteq\mathcal{T}^{\prime}$ such that $|\mathcal{T}^{\prime}_{X}|=t$, and for all $T\in\mathcal{T}^{\prime}_{X}$, we have $N_{M}(T)=X$. Observe that as $\mathcal{T}^{\prime}\subseteq\mathcal{T}^{+}$, we have $|X|\geq t$, so we can take any subset $X_{0}$ of $X$ such that $|X_{0}|=t$ and then $(X_{0},\mathcal{T}^{\prime}_{X})$ is a $\widetilde{K_{t,t}}$, as wanted.